OpenStudy (anonymous):

can someone help me make up a radical expression and convert it into a radical exponent?/????

OpenStudy (anonymous):

\[\sqrt{3^4} = 3^{\frac{4}{2}}\]

OpenStudy (anonymous):

how was the expression converted?

OpenStudy (anonymous):

you see the 3? it keeps its own exponent but takes onthe root of the radical as a bottom number...

OpenStudy (anonymous):

more...the square root of 3 is the same as 3 to the power of 1/2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

im still not uderstanding how you got it ... so your saying in the beginging before you put it in that form it was 3 sqrt 4 and waht?

OpenStudy (anonymous):

nope. to begin it was just sqrt(3^4) could have been any number really. and any radical. 5radical[16] = 16^1/5

OpenStudy (anonymous):

a radical is what happens when you get a fraction as an exponent

OpenStudy (anonymous):

so the answer is 3^4 16^1/5???

OpenStudy (anonymous):

what answer? you wanted someone to help you make radical into exponent. there is no "answer" fo be figured out. just examples.

OpenStudy (anonymous):

yes so is it 3^4 16 ^1/5

OpenStudy (anonymous):

Keep it simple: sq rt 3 is same as 3 ^ 1/2. Sometimes the number has a square inside: sq rt of 3^2 becomes 3^2*(1/2)

OpenStudy (anonymous):

no. 2 examples dont make one answer lol. give me radical and i show you exponent ok?

OpenStudy (anonymous):

ok 5 sqrt 4

OpenStudy (anonymous):

5 sqrt4 can be sqrt25*4 =sqrt100 sqrt100 = 100^1/2

OpenStudy (anonymous):

give me another one...show you more :)

OpenStudy (anonymous):

3 sqrt 4

OpenStudy (anonymous):

3sqrt4 = sqrt9*4 = sqrt36 = 36^1/2

OpenStudy (anonymous):

makes good sense yet ?

OpenStudy (anonymous):

yes thanks